# Complex exponential integral means spectrums of univalent functions and the Brennan conjecture

### v6

@@ 6.1 Univalent rational functions and the main result of this section
$R(z)=\frac{P_{n}(z)}{\prod\limits_{j=1}^{l}(z-e^{i\theta_{j}})P_{m}(z)}.$ (6.2)
@@
$R(z)=\frac{P_{n}(z)}{\prod\limits_{j=1}^{l}(z-e^{i\theta_{j}})^{2}\prod\limits_{k=1}^{t}(z-e^{i\widetilde{\theta}_{k}})P_{m}(z)}.$ (6.3)
@@
As we shall see later, the integral means spectrums of functions $R\in\mathcal{U}_{R}$ that have no critical points in $\mathbb{T}$, are different from those of functions $R$ that have at least one critical point in $\mathbb{T}$. We will list some examples of univalent rational functions $R(z)$ which has no critical points, or has at least one critical point in $\mathbb{T}$. We first consider the univalent polynomials.
@@
$R^{\prime}(z)=\frac{[(z-z_{1})(z-z_{2})\cdots(z-z_{s})]P_{\bf n}(z)}{P_{\bf m}(z)}:=\Pi(z)\frac{P_{\bf n}(z)}{P_{\bf m}(z)}.$ (6.4)
@@
We will show that the Brennan conjecture B(2)=1B(-2)=1 is true for the class $\mathcal{U}_{R}$. Actually, we shall completely determine the generalized integral means spectrums of functions belonging to the class $\mathcal{R}_{O}\cap({\mathcal{L}_{I}}\cup{\mathcal{L}_{II}}\cup{\mathcal{L}_{III}})$ for all real $\tau$. Let $\tau\in\mathbb{R}$. For each $R\in\mathcal{R}$, since $R$ has only finitely many critical points, we know that there is a constant $r_{R}\in(0,1)$ such that $R^{\prime}(z)\neq 0$ for all $z\in\{z:r_{R}<|z|<1\}$. We define the generalized integral means spectrum ${\widetilde{\beta}}_{R}$ of $R$ as
${\widetilde{\beta}}_{R}(\tau):=\lim\limits_{r\rightarrow 1^{-}}\sup\limits_{r\in(r_{R},1)}\frac{\log\int_{-\pi}^{\pi}|R^{\prime}(re^{i\theta})|^{\tau}d\theta}{|\log(1-r)|}.$ (6.7)
@@
$\int_{-\alpha}^{\alpha}\frac{d\theta}{|1-re^{i\theta}|^{\varkappa}}\asymp\frac{1}{(1-r)^{\varkappa-1}},\,\text{as}\,\,r\rightarrow 1^{-},$ (6.8)
@@
$\int_{-\alpha}^{\alpha}\frac{d\theta}{|1-re^{i\theta}|^{\varkappa}}\asymp\log\frac{1}{1-r},\,\text{as}\,\,r\rightarrow 1^{-},$ (6.9)
@@
$\int_{-\alpha}^{\alpha}\frac{d\theta}{|1-re^{i\theta}|^{\varkappa}}\asymp 1,\,\text{as}\,\,r\rightarrow 1^{-}.$ (6.10)
@@
$1-\frac{\theta^{2}}{2}\leq\cos\theta\leq 1-\frac{\theta^{2}}{4}.$ (6.11)
@@
$\int_{-\alpha}^{\alpha}\frac{d\theta}{|1-re^{i\theta}|^{\varkappa}}\asymp\frac{1}{(1-r)^{\varkappa-1}},\,\text{as}\,\,r\rightarrow 1^{-},$ (6.14)
@@
$\int_{-\alpha}^{\alpha}\frac{d\theta}{|1-re^{i\theta}|^{\varkappa}}>\int_{-\frac{\pi}{2}}^{\frac{\pi}{2}}\frac{d\theta}{|1-re^{i\theta}|^{\varkappa}}.$ (6.15)
@@
$\int_{-\alpha}^{\alpha}\frac{d\theta}{|1-re^{i\theta}|^{\varkappa}}$ $=$ $\int_{-\frac{\pi}{2}}^{\frac{\pi}{2}}\frac{d\theta}{|1-re^{i\theta}|^{\varkappa}}+\int_{\frac{\pi}{2}}^{\alpha}\frac{d\theta}{|1-re^{i\theta}|^{\varkappa}}+\int_{-\alpha}^{-\frac{\pi}{2}}\frac{d\theta}{|1-re^{i\theta}|^{\varkappa}}$ (6.16)
@@
$\int_{-\pi}^{\pi}\frac{d\theta}{|f(re^{i\theta})|^{\varkappa}}\asymp\begin{cases}\frac{1}{(1-r)^{\varkappa-1}},\;\text{if}\;\varkappa>1,\\ \log\frac{1}{1-r},\;\;\;\text{if}\;\varkappa=1,\\ 1,\quad\quad\quad\;\,\text{if}\;\varkappa\in(0,1).\end{cases}$ (6.17)
@@
$\int_{-\pi}^{\pi}\frac{d\theta}{|f(re^{i\theta})|^{\varkappa}}\geq\sum_{j=1}^{n}\int_{-{\alpha}_{j}+\arg z_{j}}^{\alpha_{j}+\arg z_{j}}\frac{d\theta}{|f(re^{i\theta})|^{\varkappa}}.$ (6.18)
@@
$\int_{-\pi}^{\pi}\frac{d\theta}{|R^{\prime}(re^{i\theta})|^{-\tau}}\asymp\begin{cases}\frac{1}{(1-r)^{-\tau-1}},\;\;\text{if}\;\tau<-1,\\ \log\frac{1}{1-r},\;\;\;\;\;\text{if}\;\tau=-1,\\ 1,\quad\quad\quad\;\,\;\;\text{if}\;\tau\in(-1,0).\end{cases}$ (6.22)
@@
${\widetilde{\beta}}_{R}(\tau)=\begin{cases}3\tau-1,\,\text{for}\;\tau>\frac{1}{3},\\ 0,\quad\,\,\;\;\;\;\text{for}\;\tau\in(0,\frac{1}{3}].\end{cases}$ (6.23)
@@
7 Some remarks
7.1 The first part of the remarks
@@
7.2 The second part of the remarks
@@
$\kappa^{\prime}_{T}(z)=\frac{(1+z^{T})}{(1-z^{T})^{1+\frac{2}{T}}},z\in\Delta,$ (7.2)
@@
In particular, $S_{\kappa_{2}}(z)=\frac{6}{(1+z^{2})^{2}}.$ In view of Lemma 6.24, we obtain from (7.2) that $\beta_{\kappa_{T}}(\tau)=|\tau|-1$ for each $T\in\mathbb{N}$ when $\tau\leq-2$. Also, it is not hard to check that, for each $T\in\mathbb{N}$,
@@
7.3 The third part of the remarks
@@
$\lim\limits_{\gamma\rightarrow 1^{-}}\|N_{g_{\gamma}}-N_{\mathcal{E}_{2}}\|_{E_{1}}=0.$ (7.3)
@@
7.4 The final part of the remarks
In the paper [52], Shimorin introduced and studied the following multiplier operator $\mathcal{M}_{f}$, induced by the Schwarzian derivative of univalent function $f\in\mathcal{S}$,
@@ 7.4 The final part of the remarks
There is a close connection between the norm of multiplier operator $\mathcal{M}_{f}$ and the Brennan conjecture. It has been proved in [52] that
@@
$\|\mathcal{M}_{f}\|^{2}\leq\frac{36(\alpha+3)(\alpha+5)}{(\alpha+2)(\alpha+4)},$ (7.4)
@@
It is easy to see that, if for any $\alpha>0$, (7.4) holds for all $f\in\mathcal{S}_{q}$, then the Brennan conjecture is true. On the other hand, for any $\alpha>-1$, it has been shown in [52] that,
$\mathbb{M}_{\kappa_{\Theta}}^{2}(\alpha)=\frac{36(\alpha+3)(\alpha+5)}{(\alpha+2)(\alpha+4)}.$ (7.5)
@@
$\kappa_{\Theta}(z):=\frac{z}{(1-e^{i\Theta}z)^{2}},\,\Theta\in(-\pi,\pi],\,z\in\Delta.$ (7.6)
@@
Since the Schwarzian derivative is invariant under the Möbius transformation, we see that, for any $\alpha>-1$, it holds that $\mathbb{M}_{\varrho_{c}\circ\kappa_{\Theta}}(\alpha)=\mathbb{M}_{\kappa}(\alpha),$ for any $(c,\Theta)\in[-4,4]\times(-\pi,\pi]$. Because the Koebe function and its rotations are the unique class of extremal functions for many extremal problems in the theory of univalent functions, thus, it is natural to guess that the functions $\varrho_{c}\circ\kappa_{\Theta}$, $(c,\Theta)\in[-4,4]\times(-\pi,\pi]$, may induce the biggest norm of the operator $\mathcal{M}_{f}$ in the class $\mathcal{S}$ for any $\alpha>0$, However, we find that there are other univalent functions $f\in\mathcal{S}$ such that (7.5) still holds for any $\alpha>-1$. Meanwhile, it should be pointed out that there exists univalent function $f\in\mathcal{S}$ such that $\mathbb{M}_{f}(\alpha)>\mathbb{M}_{\kappa}(\alpha)$ when $\alpha>-1$, see [52].
Now, we consider again the functions $\mathcal{E}_{2}$ and $P_{3}$. We will check that
@@
$\frac{1}{(1-z)^{\gamma}}=\sum_{n=0}^{\infty}\frac{\Gamma(n+\gamma)}{n!\Gamma(\gamma)}z^{n},z\in\Delta.$ (7.7)
@@
$\|\phi\|_{\alpha}^{2}=\sum_{n=0}^{\infty}\frac{n!\Gamma(\alpha+2)}{\Gamma(n+\alpha+2)}|a_{n}|^{2}.$ (7.8)
@@
$\|{(1-rz)^{-\lambda}}\|_{\alpha}^{2}=\sum_{n=0}^{\infty}\frac{n!\Gamma(\alpha+2)}{\Gamma(n+\alpha+2)}\Big|\frac{\Gamma(n+\lambda)}{n!\Gamma(\lambda)}\Big|^{2}r^{2n}.$ (7.9)
@@
$\|{(1-rz)^{-\lambda}}\|_{\alpha}^{2}$ $=$ $\sum_{n=0}^{\infty}\frac{\Gamma(\alpha+2)}{n^{\alpha+1}}[1+o(1)]\cdot\frac{n^{2\lambda-2}}{[\Gamma(\lambda)]^{2}}[1+o(1)]r^{2n}$ (7.10)
@@
$[{(1-r^{2})^{2\lambda-\alpha-2}}]\sum_{n=0}^{\infty}\frac{\Gamma(n+2\lambda-\alpha-2)}{n!\Gamma(2\lambda-\alpha-2)}[1+o(1)]r^{2n}$ (7.11)
@@
$[{(1-r^{2})^{2\lambda-\alpha-2}}]\sum_{n=0}^{\infty}\frac{\Gamma(n+2\lambda-\alpha-2)}{n!\Gamma(2\lambda-\alpha-2)}\cdot[o(1)]\cdot r^{2n}$ (7.12)
@@
Can the functional $\Lambda_{\alpha}$ attain at least one maximum on $\overline{\mathbf{S}}_{Q}$ for each $\alpha>-1$?
@@
In light of Proposition 7.26, it is interesting to study the problem that in which subset of $\mathcal{S}_{Q}$ the Koebe function and its rotations provide the biggest norm for the multiplier operator when $\alpha>0$. That is
@@
For every $\alpha>0$, which functions $f\in\mathcal{S}_{Q}$ satisfy $\mathbb{M}_{f}(\alpha)\leq\mathbb{M}_{\kappa}(\alpha)$?
@@
for $\alpha\in(0,\varepsilon)$, then $\beta_{f}({-2})\leq 1$. Moreover, for $R\in\mathcal{U}_{R}$ that have at least one critical point in $\mathbb{T}$, if there are two small constants $\epsilon>0,\eta>0$ such that $\mathbb{M}_{R}(\alpha)<\mathbb{M}_{\kappa}(\alpha)-\eta$ for $\alpha\in(-\epsilon,\epsilon)\cup(2-\epsilon,2+\epsilon)$, then we will obtain that $\beta_{R}(-2)<1$, which is a contradiction. From this, we believe that the following problem has a positive answer.
Problem 7.37.
For every $\alpha>0$, does it hold that $\mathbb{M}_{R}(\alpha)\geq\mathbb{M}_{\kappa}(\alpha)$ for all $R\in\mathcal{U}_{R}$ that have at least one critical point in $\mathbb{T}$?
We conclude by remarking that, if Problem 7.35 has an affirmative answer for every $f\in\mathcal{S}_{Q}$, i.e., if $\mathbb{M}f(\alpha)\leq\mathbb{M}{\kappa}(\alpha)$ holds for all $\alpha>0$ and all $f\in\mathcal{S}{Q}$, then the Brennan conjecture follows. Furthermore, if Problem 7.37 also has an affirmative answer, then it follows that $\mathbb{M}_{R}(\alpha)=\mathbb{M}_{\kappa}(\alpha)$ for every $\alpha>0$ and all $R\in\mathcal{U}_{R}$ that have at least one critical point in $\mathbb{T}$. While these predictions might appear surprising, no counterexamples are known to the authors.

### v7

@@ 6.1. Univalent rational functions and the main result of this section
(6.2) $R(z)=\frac{P_{n}(z)}{\prod\limits_{j=1}^{l}(z-e^{i\theta_{j}})P_{m}(z)}.$
@@
(6.3) $R(z)=\frac{P_{n}(z)}{\prod\limits_{j=1}^{l}(z-e^{i\theta_{j}})^{2}\prod\limits_{k=1}^{t}(z-e^{i\widetilde{\theta}_{k}})P_{m}(z)}.$
@@
As we shall see later, the integral means spectra of functions $R\in\mathcal{U}_{R}$ that have no critical points in $\mathbb{T}$, are different from those of functions $R$ that have at least one critical point in $\mathbb{T}$. We will list some examples of univalent rational functions $R(z)$ which has no critical points, or has at least one critical point in $\mathbb{T}$. We first consider the univalent polynomials.
@@
(6.4) $R^{\prime}(z)=\frac{[(z-z_{1})(z-z_{2})\cdots(z-z_{s})]P_{\bf n}(z)}{P_{\bf m}(z)}:=\Pi(z)\frac{P_{\bf n}(z)}{P_{\bf m}(z)}.$
@@
We will show that the Brennan conjecture B(2)=1B(-2)=1 is true for the class $\mathcal{U}_{R}$. Actually, we shall completely determine the generalized integral means spectra of functions belonging to the class $\mathcal{R}_{O}\cap({\mathcal{L}_{I}}\cup{\mathcal{L}_{II}}\cup{\mathcal{L}_{III}})$ for all real $\tau$. Let $\tau\in\mathbb{R}$. For each $R\in\mathcal{R}$, since $R$ has only finitely many critical points, we know that there is a constant $r_{R}\in(0,1)$ such that $R^{\prime}(z)\neq 0$ for all $z\in\{z:r_{R}<|z|<1\}$. We define the generalized integral means spectrum ${\widetilde{\beta}}_{R}$ of $R$ as
(6.7) ${\widetilde{\beta}}_{R}(\tau):=\lim\limits_{r\rightarrow 1^{-}}\sup\limits_{r\in(r_{R},1)}\frac{\log\int_{-\pi}^{\pi}|R^{\prime}(re^{i\theta})|^{\tau}d\theta}{|\log(1-r)|}.$
@@
(6.8) $\int_{-\alpha}^{\alpha}\frac{d\theta}{|1-re^{i\theta}|^{\varkappa}}\asymp\frac{1}{(1-r)^{\varkappa-1}},\,\text{as}\,\,r\rightarrow 1^{-},$
@@
(6.9) $\int_{-\alpha}^{\alpha}\frac{d\theta}{|1-re^{i\theta}|^{\varkappa}}\asymp\log\frac{1}{1-r},\,\text{as}\,\,r\rightarrow 1^{-},$
@@
(6.10) $\int_{-\alpha}^{\alpha}\frac{d\theta}{|1-re^{i\theta}|^{\varkappa}}\asymp 1,\,\text{as}\,\,r\rightarrow 1^{-}.$
@@
(6.11) $1-\frac{\theta^{2}}{2}\leq\cos\theta\leq 1-\frac{\theta^{2}}{4}.$
@@
(6.14) $\int_{-\alpha}^{\alpha}\frac{d\theta}{|1-re^{i\theta}|^{\varkappa}}\asymp\frac{1}{(1-r)^{\varkappa-1}},\,\text{as}\,\,r\rightarrow 1^{-},$
@@
(6.15) $\int_{-\alpha}^{\alpha}\frac{d\theta}{|1-re^{i\theta}|^{\varkappa}}>\int_{-\frac{\pi}{2}}^{\frac{\pi}{2}}\frac{d\theta}{|1-re^{i\theta}|^{\varkappa}}.$
@@
(6.16) $\int_{-\alpha}^{\alpha}\frac{d\theta}{|1-re^{i\theta}|^{\varkappa}}$ $=$ $\int_{-\frac{\pi}{2}}^{\frac{\pi}{2}}\frac{d\theta}{|1-re^{i\theta}|^{\varkappa}}+\int_{\frac{\pi}{2}}^{\alpha}\frac{d\theta}{|1-re^{i\theta}|^{\varkappa}}+\int_{-\alpha}^{-\frac{\pi}{2}}\frac{d\theta}{|1-re^{i\theta}|^{\varkappa}}$
@@
(6.17) $\int_{-\pi}^{\pi}\frac{d\theta}{|f(re^{i\theta})|^{\varkappa}}\asymp\begin{cases}\frac{1}{(1-r)^{\varkappa-1}},\;\text{if}\;\varkappa>1,\\ \log\frac{1}{1-r},\;\;\;\text{if}\;\varkappa=1,\\ 1,\quad\quad\quad\;\,\text{if}\;\varkappa\in(0,1).\end{cases}$
@@
(6.18) $\int_{-\pi}^{\pi}\frac{d\theta}{|f(re^{i\theta})|^{\varkappa}}\geq\sum_{j=1}^{n}\int_{-{\alpha}_{j}+\arg z_{j}}^{\alpha_{j}+\arg z_{j}}\frac{d\theta}{|f(re^{i\theta})|^{\varkappa}}.$
@@
(6.22) $\int_{-\pi}^{\pi}\frac{d\theta}{|R^{\prime}(re^{i\theta})|^{-\tau}}\asymp\begin{cases}\frac{1}{(1-r)^{-\tau-1}},\;\;\text{if}\;\tau<-1,\\ \log\frac{1}{1-r},\;\;\;\;\;\text{if}\;\tau=-1,\\ 1,\quad\quad\quad\;\,\;\;\text{if}\;\tau\in(-1,0).\end{cases}$
@@
(6.23) ${\widetilde{\beta}}_{R}(\tau)=\begin{cases}3\tau-1,\,\text{for}\;\tau>\frac{1}{3},\\ 0,\quad\,\,\;\;\;\;\text{for}\;\tau\in(0,\frac{1}{3}].\end{cases}$
@@
7. Some remarks
7.1. The first part of the remarks
@@
7.2. The second part of the remarks
@@
(7.2) $\kappa^{\prime}_{T}(z)=\frac{(1+z^{T})}{(1-z^{T})^{1+\frac{2}{T}}},z\in\Delta,$
@@
In particular,
$S_{\kappa_{2}}(z)=\frac{6}{(1+z^{2})^{2}}.$
In view of Lemma 6.24, we obtain from (7.2) that $\beta_{\kappa_{T}}(\tau)=|\tau|-1$ for each $T\in\mathbb{N}$ when $\tau\leq-2$. Also, it is not hard to check that, for each $T\in\mathbb{N}$,
@@
7.3. The third part of the remarks
@@
(7.3) $\lim\limits_{\gamma\rightarrow 1^{-}}\|N_{g_{\gamma}}-N_{\mathcal{E}_{2}}\|_{E_{1}}=0.$
@@
7.4. The final part of the remarks
In the paper [52], Shimorin introduced and studied the following multiplication operator $\mathcal{M}_{f}$, induced by the Schwarzian derivative of univalent function $f\in\mathcal{S}$,
@@ 7.4. The final part of the remarks
There is a close connection between the norm of multiplication operator $\mathcal{M}_{f}$ and the Brennan conjecture. It has been proved in [52] that
@@
(7.4) $\|\mathcal{M}_{f}\|^{2}\leq\frac{36(\alpha+3)(\alpha+5)}{(\alpha+2)(\alpha+4)},$
@@
It is easy to see that, if for any $\alpha>0$, (7.4) holds for all $f\in\mathcal{S}_{q}$, then the Brennan conjecture is true. On the other hand, it has been pointed out in [52, Page 1633] that there exists univalent function $f\in\mathcal{S}$ such that $\mathbb{M}_{f}(\alpha)>\mathbb{M}_{\kappa}(\alpha)$ when $\alpha>-1$. Meanwhile, for any $\alpha>-1$, it has been shown in [52] that,
(7.5) $\mathbb{M}_{\kappa_{\Theta}}^{2}(\alpha)=\frac{36(\alpha+3)(\alpha+5)}{(\alpha+2)(\alpha+4)}.$
@@
(7.6) $\kappa_{\Theta}(z):=\frac{z}{(1-e^{i\Theta}z)^{2}},\,\Theta\in(-\pi,\pi],\,z\in\Delta.$
@@
Since the Schwarzian derivative is invariant under the Möbius transformation, we see that, for any $\alpha>-1$, it holds that $\mathbb{M}_{\varrho_{c}\circ\kappa_{\Theta}}(\alpha)=\mathbb{M}_{\kappa}(\alpha),$ for any $(c,\Theta)\in[-4,4]\times(-\pi,\pi]$. Because the Koebe function and its rotations are the unique class of extremal functions for some extremal problems in the theory of univalent functions, thus, it is natural to guess that the functions $\varrho_{c}\circ\kappa_{\Theta}$, $(c,\Theta)\in[-4,4]\times(-\pi,\pi]$, may induce the biggest norm of the operator $\mathcal{M}_{f}$ in the class $\mathcal{S}$ for any $\alpha>0$. However, we find that there are other univalent functions $f\in\mathcal{S}$ such that (7.5) still holds for any $\alpha>-1$.
First, we note that for any $\theta\in\mathbb{R}$, $S_{\kappa}(e^{i\theta}z)$ induces the same multiplication norm as $S_{\kappa}(z)$(see [52, Page 1631]), then we see from Example 7.12 that $\mathbb{M}_{\kappa_{2}}(\alpha)=\mathbb{M}_{\kappa}(\alpha)$ for each $\alpha>-1$. Next, we consider again the functions $\mathcal{E}_{2}$ and $P_{3}$. We will check that
@@
(7.7) $\frac{1}{(1-z)^{\gamma}}=\sum_{n=0}^{\infty}\frac{\Gamma(n+\gamma)}{n!\Gamma(\gamma)}z^{n},z\in\Delta.$
@@
(7.8) $\|\phi\|_{\alpha}^{2}=\sum_{n=0}^{\infty}\frac{n!\Gamma(\alpha+2)}{\Gamma(n+\alpha+2)}|a_{n}|^{2}.$
@@
(7.9) $\|{(1-rz)^{-\lambda}}\|_{\alpha}^{2}=\sum_{n=0}^{\infty}\frac{n!\Gamma(\alpha+2)}{\Gamma(n+\alpha+2)}\Big|\frac{\Gamma(n+\lambda)}{n!\Gamma(\lambda)}\Big|^{2}r^{2n}.$
@@
(7.10) $\|{(1-rz)^{-\lambda}}\|_{\alpha}^{2}$ $=$ $\sum_{n=0}^{\infty}\frac{\Gamma(\alpha+2)}{n^{\alpha+1}}[1+o(1)]\cdot\frac{n^{2\lambda-2}}{[\Gamma(\lambda)]^{2}}[1+o(1)]r^{2n}$
@@
(7.11) $[{(1-r^{2})^{2\lambda-\alpha-2}}]\sum_{n=0}^{\infty}\frac{\Gamma(n+2\lambda-\alpha-2)}{n!\Gamma(2\lambda-\alpha-2)}[1+o(1)]r^{2n}$
@@
(7.12) $[{(1-r^{2})^{2\lambda-\alpha-2}}]\sum_{n=0}^{\infty}\frac{\Gamma(n+2\lambda-\alpha-2)}{n!\Gamma(2\lambda-\alpha-2)}\cdot[o(1)]\cdot r^{2n}$
@@
Can the functional $\Lambda_{\alpha}$ attain at least one maximum in $\overline{\mathbf{S}}_{Q}$ for each $\alpha>-1$?
@@
In light of Proposition 7.26, it is interesting to study the problem that in which subset of $\mathcal{S}_{Q}$ the Koebe function and its rotations provide the biggest norm for the multiplication operator when $\alpha>0$. That is
@@
For each $\alpha>0$, which functions $f\in\mathcal{S}_{Q}$ satisfy $\mathbb{M}_{f}(\alpha)\leq\mathbb{M}_{\kappa}(\alpha)$?
@@
for $\alpha\in(0,\varepsilon)$, then $\beta_{f}({-2})\leq 1$. Moreover, for $R\in\mathcal{U}_{R}$ that have at least one critical point in $\mathbb{T}$, if there are two small constants $\epsilon>0,\eta>0$ such that $\mathbb{M}_{R}(\alpha)<\mathbb{M}_{\kappa}(\alpha)-\eta$ for $\alpha\in(-\epsilon,\epsilon)\cup(2-\epsilon,2+\epsilon)$, then we will obtain that $\beta_{R}(-2)<1$, which is a contradiction. This leads us to conjecture that the following claim holds. Indeed, this is the case, and a proof will be given elsewhere.
Claim 7.37.
For each $\alpha>-1$, we have $\mathbb{M}_{R}(\alpha)\geq\mathbb{M}_{\kappa}(\alpha)$ for all $R\in\mathcal{U}_{R}$ that have at least one critical point in $\mathbb{T}$.
Remark 7.38.
We have proved that, for each $\alpha>-1$,
$\mathbb{M}_{\mathcal{E}_{2}}^{2}(\alpha)=\mathbb{M}_{P_{3}}^{2}(\alpha)=\mathbb{M}_{\kappa}^{2}(\alpha)=\frac{36(\alpha+3)(\alpha+5)}{(\alpha+2)(\alpha+4)}.$
It is natural to consider the following
Problem 7.39.
For each $\alpha>-1$, which functions $f\in\mathcal{U}_{R}$ satisfy $\mathbb{M}_{f}(\alpha)=\mathbb{M}_{\kappa}(\alpha)$?
Remark 7.40.
We conclude by remarking that, if our expectation that Problem 7.35 has an affirmative answer for every $f\in\mathcal{S}_{Q}$ is valid, i.e., if $\mathbb{M}f(\alpha)\leq\mathbb{M}{\kappa}(\alpha)$ holds for all $\alpha>0$ and all $f\in\mathcal{S}_{Q}$, then the Brennan conjecture follows. Furthermore, since Claim 7.37 holds, then it follows that $\mathbb{M}_{R}(\alpha)=\mathbb{M}_{\kappa}(\alpha)$ for every $\alpha>0$ and all $R\in\mathcal{U}_{R}$ that have at least one critical point in $\mathbb{T}$. Although these predictions may seem surprising, no counterexamples are known to the authors.